OpenStudy (anonymous):

state the domain. 6/4x^3

jimthompson5910 (jim_thompson5910):

is the expression \[\large \frac{6}{4}x^3\] or is it \[\large \frac{6}{4x^3}\]

OpenStudy (anonymous):

the secondone

jimthompson5910 (jim_thompson5910):

since you cannot divide by zero, the denominator 4x^3 cannot be zero

jimthompson5910 (jim_thompson5910):

so 4x^3 = 0 leads to x = ??